# The building blocks of software work explain coding careers and language popularity

### v1

@@ Abstract
Recent waves of technological transformation have fueled debates about the changing nature of work. Yet to understand the future of work, we need to know more about what people actually do in their jobs, going beyond educational credentials or job descriptions. Here we analyze work in the global software industry using tens of millions of Question and Answer posts on Stack Overflow to create a fine-grained taxonomy of software tasks, the elementary building blocks of software development work. These tasks predict salaries and job requirements in real-world job ads. We also observe how individuals learn within tasks and diversify into new tasks. Tasks that people acquire tend to be related to their old ones, but of lower value, suggesting that they are easier. An exception is users of Python, an increasingly popular programming language known for its versatility. Python users enter tasks that tend to be higher-value, providing an explanation for the language’s growing popularity based on the tasks Python enables its users to perform. In general, these insights demonstrate the value of task taxonomies extracted at scale from large datasets: they offer high resolution and near real-time descriptions of changing labor markets. In the case of software tasks, they map such changes for jobs at the forefront of a digitizing global economy.
Classification: Social Sciences/Economics
Keywords: human capital; tasks; software; economic complexity; networks
@@ 1 Introduction
Knowledge of the tasks that workers perform is central to understanding wages, productivity, and overall labor market dynamics (Acemoglu and Autor, 2011). Recent work has transformed our ability to do so by using complex network analysis — moving away from descriptions that rely on years of schooling, credentials, and work experience to the granular building blocks of human capital: skills and tasks (Anderson, 2017; Alabdulkareem et al., 2018; Hosseinioun et al., 2025). These studies focus on jobs, describing them as networks of skills or tasks using data on skill requirements and task descriptions extracted from large-scale surveys (one, 2025), expert assessments (Commission, 2017), job ads (Börner et al., 2018) or online freelance platforms(Anderson, 2017; Stephany and Teutloff, 2024). Yet, job descriptions tell us something about the nature of vacancies, not of the people who fulfill them. This limits our ability to understand how workers specialize and learn new things, processes critical to a successful career.
Focusing on the software sector, we propose a new approach that uses data from Stack Overflow (SO), a very large, online question-answer database. The core insight we leverage is that questions on SO describe coding problems that programmers encounter in their work. As such, they can be understood as instantiations of software-development job-tasks. By generalizing these descriptions of specific problems encountered at work, we can construct a taxonomy of canonical software tasks that is highly granular, flexible and based in data. Moreover, we can use this taxonomy to describe the human capital endowments of people, not the human capital requirements of jobs. That is, we ask what programmers do at work, not what they are supposed to do according to their job descriptions. This allows us to analyze how programmers learn new tasks and what the value of these tasks is. Moreover, programmers can choose from a wide menu of programming languages and we show that programming languages differ in the tasks they support. However, the mapping of tasks to programming languages changes over time, as illustrated by the rapid rise of Python as a general purpose language. Our task network shows that this rise exploits a remarkable strength of Python: it allows programmers to more quickly transition to the most valuable programming tasks.
As wages and careers have come to depend on how workers specialize (Hosseinioun et al., 2025), interact (Deming, 2017), complement colleagues (Neffke, 2019), and move across places and jobs (Frank et al., 2024), fine-grained taxonomies of tasks are increasingly valuable. Such taxonomies clarify how workers transition between occupations (del Rio-Chanona et al., 2021), how skills can substitute for or complement each other (Anderson et al., 2012; Neffke, 2019), and how occupations evolve over time (Nedelkoska et al., 2021).
@@
In this paper, we focus on jobs in the global software development sector. Software production is a fast-changing and critical domain of the digital economy (Brynjolfsson and McAfee, 2014; Wachs et al., 2022; Juhász et al., 2024). Of the seven largest companies by market capitalization worldwide, three (Microsoft, Alphabet, and Meta) focus almost exclusively on software, while three others (Nvidia, Apple, and Amazon) derive substantial revenue from it. Demand for programming skills continues to surge: one in 11 U.S. occupations now requires coding (one, 2025), and software-related roles such as data scientist, information security analyst, and computer research scientist rank among the fastest-growing occupations (U.S. Bureau of Labor Statistics, 2025). As new hardware, programming languages, and specialized libraries reshape computing, fresh roles (e.g., web development, DevOps, data science, artificial intelligence) keep emerging (Rock, 2019; Papoutsoglou et al., 2019; Hemon et al., 2020). This changing landscape has sparked intense global competition for talent, prompting many major economies to designate software development as a strategic priority (Draghi, 2024; Kazim et al., 2021; Singh, 2024).
@@
However, we currently lack a detailed understanding of the fast-changing human capital that is employed in this sector. One reason is that databases linking occupations to tasks (e.g., O*NET (one, 2025)) are often curated by human experts, making them relatively coarse and slow to update. In O*NET, for example, software-related tasks often focus on broad domains of application (e.g., “geoinformatics” or “biostatistics”), team dynamics (e.g., managing or collaborating with software engineers), or acquiring software products. Alternative data sources such as job advertisements or resumes are also challenging because job ads describe job requirements of ideal-type workers, not human capital endowments of actual workers, resumes can be aspirational, and both often focus on limited geographies. Likewise, platforms for online freelancing provide insights into how complex skill interdependencies affect wage premiums (Anderson, 2017; Stephany and Teutloff, 2024), but the extent to which these findings generalize to the broader labor market remains uncertain.
We address these gaps by analyzing millions of posts from SO—the largest question-and-answer forum for programmers (Anderson et al., 2012). Each post describes a specific technical challenge and includes tags that indicate the tools, languages, or concepts involved (Barua et al., 2014). These challenges reflect real-world coding tasks and users use their posting histories in part to signal their skills to prospective employers (Xu et al., 2020). Interpreting tags as a structured, albeit partial, description of posted questions, we aggregate questions to create a software-task taxonomy by identifying clusters of frequently co-occurring tags. Using this taxonomy, we examine how software tasks evolve in the labor market. We find that task value estimates, derived from survey data, significantly predict salaries in real-world job postings. Notably, wages vary greatly among these tasks, with higher-paying ones concentrated in more specialized or in-demand fields such as AI and machine learning. We also introduce a notion of task relatedness based on co-occurrence patterns in users’ posting histories. This measure of relatedness derived from interactions with the SO platform predicts similar co-occurrence trends in job ads, suggesting that our taxonomy captures a meaningful structure of the software development landscape.
At the individual user level, we observe two learning patterns. First, developers improve in the tasks they gain experience in, as measured by community feedback on their posts. Second, developers enter new tasks over their SO careers, and task relatedness strongly influences these entries: developers are fifteen times more likely to move into tasks related to ones they already know. At the same time, we find that developers usually add tasks with lower market value, which are presumably easier to master.
@@
Finally we demonstrate that our fine-grained software task taxonomy sheds light on the evolution of software development as a whole. We observe that programming languages differ in the number of tasks for which they are used, following a nested pattern of tasks in programming languages reminiscent of similar patterns observed in ecology (Bascompte et al., 2003), economic development (Hidalgo and Hausmann, 2009; Tacchella et al., 2012; Mariani et al., 2019) and skill hierarchies (Hosseinioun et al., 2025). Accordingly, instead of specialization, where different languages are used to perform different types of tasks, we observe that tasks have an implicit hierarchy such that the most ubiquitous tasks can be performed in all languages, whereas rare tasks are only performed by the most versatile ones. A prime example of the latter is the Python programming language, a popular language that has seen significant growth in recent years (Economist, 2018). We show that this growth has been accompanied by changes in the way Python is used. First, Python has become increasingly versatile. From 2015, the number of tasks in which Python dominates as the preferred language for most programmers has rapidly grown, rendering Python into one of today’s premier general purpose languages. Second, Python enables qualitatively different paths for users to learn new tasks. Specifically, unlike users of most other languages, Python users often branch into more, not less valuable tasks. This observation could help explain Python’s success, despite its not being the oldest or fastest language: its ecosystem may make it easier for developers to tackle higher-value tasks.

### v2

# Using digital traces to analyze software work: skills, careers and programming languages

@@ Abstract
Recent waves of technological transformation are reshaping work in uncertain and hard-to-predict ways. However, jobs at the forefront of the digitizing economy offer an early glimpse of these changes and leave rich activity traces. We exploit such traces in tens of millions of Question and Answer posts on Stack Overflow for the creation of a fine-grained taxonomy of software skills to analyze human capital in the global software industry. Constructing a software skill space that maps relations among these skills reveals that real-world software jobs demand highly coherent skill sets and that programmers learn through a process of related diversification. The latter process often leads to the acquisition of lower-value skills. However, when programmers use Python they preferentially target higher-value skills, offering a potential explanation for Python’s successful rise as a dominant general purpose language.
Keywords: human capital; skills; software; economic complexity; networks
@@ 1 Introduction
Knowledge of detailed skills and expertise used at work is central to understanding wages, productivity, and overall labor market dynamics (Acemoglu and Autor, 2011). Recent work has transformed our ability to do so by using complex network analysis. This has allowed moving away from descriptions that rely on years of schooling, credentials, and work experience to studies of the relations among concrete skills and the tasks they enable workers to perform (Anderson, 2017; Alabdulkareem et al., 2018; Hosseinioun et al., 2025). However, most of these studies focus on the content of jobs, not the human capital of who performs them. This limits our understanding of how workers learn new skills, which is crucial for helping individuals, firms and educators navigate transforming labor markets in times of rapid technological change. To address this, we focus on a specific type of work that has become increasingly important, is changing fast and can be studied in great detail: software development. Constructing a fine-grained taxonomy of programming skills, we create a software development skill space that maps relations among these skills. We use this space to analyze which skills are combined in job ads, how programmers acquire new skills and what role programming languages play therein.
@@
Software jobs are particularly interesting, because they represent a fast-changing and critical domain of the digital economy (Brynjolfsson and McAfee, 2014; Wachs et al., 2022; Juhász et al., 2026). Software companies dominate the global ranks of the largest firms by market capitalization. [^1] Of the seven largest companies worldwide, three (Microsoft, Alphabet, and Meta) focus almost exclusively on software, while two others (Apple and Amazon) derive substantial revenue from it. On the labor market, demand for programming skills has risen sharply over the past two decades: one in 11 U.S. occupations now requires coding skills (O*NET, 2025), and software-related roles such as data scientist, information security analyst, and computer research scientist rank among the fastest-growing occupations in the US (U.S. Bureau of Labor Statistics, 2025). At the macro-economic level, this has sparked intense global competition for talent, prompting major economies to designate software development as a strategic priority (Draghi, 2024; Kazim et al., 2021; Singh, 2024).
@@
However, we currently lack a detailed understanding of the changing human capital that is employed in this sector. As new hardware, programming languages, and specialized libraries reshape computing, fresh roles (e.g., web development, DevOps, data science, artificial intelligence) keep emerging (Rock, 2019; Papoutsoglou et al., 2019; Hemon et al., 2020) and existing roles transform. Existing databases linking occupations to skills (e.g., O*NET, 2025)) are often curated by human experts, making them relatively coarse and slow to update. [^2] For example, the only software-related skill in O*NET is undifferentiated “programming” and the listed software-related work activities often focus on broad domains of application (e.g., “geoinformatics” or “biostatistics”), team dynamics (e.g., managing or collaborating with software engineers), or acquiring software products. More dynamic data sources are challenging in their own right. For instance, job ads — frequently used to study job contents — describe job requirements of ideal-type workers, not human capital endowments of actual workers. Likewise, although rich data from platforms that mediate online freelancing work provide insights into how complex skill interdependencies affect wage premiums (Anderson, 2017; Stephany and Teutloff, 2024), the extent to which these findings generalize to traditional labor markets remains uncertain.
We address these gaps by analyzing millions of posts from Stack Overflow (SO)—the largest question-and-answer forum for programmers (Anderson et al., 2012). Each post describes a specific technical challenge and includes tags that indicate the tools, languages, or concepts involved (Barua et al., 2014). These challenges reflect real-world coding tasks and by posting answers users demonstrate skills and expertise in these tasks. Indeed, SO postings have real-world value as signals to prospective employers of a programmer’s capabilities (Xu et al., 2020).
Interpreting tags as a structured, albeit partial, description of posted questions, we aggregate questions to create a software-skill taxonomy by identifying clusters of frequently co-occurring tags. We then introduce a notion of relatedness among programming skills based on co-occurrence patterns in users’ posting histories. We use this relatedness to construct a skill space where skills are connected if they are often jointly possessed by the same users. We then use this taxonomy and skill space to examine the postings and learning trajectories of programmers.
We find that our skill constructs can be used to predict salaries in real-world job postings. Notably, wages vary greatly across skill profiles, with higher-paying jobs concentrated in more specialized or in-demand fields such as AI and machine learning. Similarly, relatedness estimates derived from interactions with the SO platform predict which combinations of skills are demanded in real-world job ads. This not only shows that our taxonomy captures meaningful structures in the software development landscape, but also that software development jobs are coherent in the sense that they often require sets of closely related skills.
@@
The skill taxonomy also sheds light on the role and evolution of programming languages. Programming languages differ in the expertise they support. Analyzing which skills are used with which programming languages, we observe a “nested” pattern reminiscent of similar patterns observed in ecology (Bascompte et al., 2003), economic development (Hidalgo and Hausmann, 2009; Tacchella et al., 2012; Mariani et al., 2019) and skill hierarchies (Hosseinioun et al., 2025). This means that although programming languages may have been tailored for specific use cases, this is not observed in how programmers use languages in practice. Instead, we observe that skills have an implicit hierarchy such that the most ubiquitous skills are supported by almost all languages, whereas rare skills often rely on the most versatile ones. A prime example of the latter is the programming language Python, a popular language that has seen significant growth in recent years (Economist, 2018). Our skill taxonomy shows that this growth has been accompanied by changes in the way Python is used, with Python becoming increasingly versatile, rapidly increasing the variety of skills it supports to become today’s premier general purpose language.
At the individual user level, we observe two patterns. First, developers excel —as inferred from community feedback on their posts— at the skills they are most active in. Second, developers acquire new skills over their SO careers, and the relatedness network strongly influences which skills they learn: developers are fifteen times more likely to acquire skills related to skills they already master than to unrelated ones. At the same time, we find that developers usually add skills with lower market value, which are presumably easier to learn. However, Python enables qualitatively different learning paths. Specifically, unlike users of most other languages, Python users often branch into more, not less valuable skills. This observation could help explain Python’s success, despite its not being the oldest or fastest language: its ecosystem may make it easier for developers to develop higher-value expertise.
Our analysis contributes to the growing literature that analyzes the content of work. This literature has shown that as wages and careers have come to depend on how workers specialize (Hosseinioun et al., 2025), interact (Deming, 2017), complement one another (Neffke, 2019), and move across places and jobs (Frank et al., 2024), fine-grained skill taxonomies are increasingly valuable. Such taxonomies clarify how workers transition between occupations (del Rio-Chanona et al., 2021), how skills substitute for or complement each other (Anderson et al., 2012; Neffke, 2019), and how occupations evolve over time (Nedelkoska et al., 2021). This line of research uses data on skill requirements and work activities extracted from large-scale surveys (O*NET, 2025), expert assessments (Commission, 2017), job ads (Börner et al., 2018) or online freelance platforms(Anderson, 2017; Stephany and Teutloff, 2024) to describe jobs as networks of skills or tasks. We contribute to this debate by focusing away from the task requirements listed in descriptions of jobs toward the human capital and skills of the people who fulfill them. This allows us to study how workers specialize and learn new things, processes critical to a successful career.
2 Mapping software work
In this section, we describe how we construct a software skill taxonomy and a space of software skills. First we describe the raw inputs: posts on Stack Overflow and their tags, labels which capture post content. We then explain how we use a network-based clustering algorithm, the stochastic block model (SBM, see Appendix A), to identify clusters of tags which frequently co-occur in posts. These clusters are our skills, from which we can derive descriptions of user skills and skill values. Finally we demonstrate how skills can be embedded in an abstract space, in which related skills are near one another.
2.1 Data
Our main dataset is the corpus of questions and answers (together “posts”) on Stack Overflow from the site’s launch in August 2008 to June 2023. The 23 million questions in this dataset are labeled by tags from a curated system of about 66,000 tags referring to software concepts and technologies such as key-value, scipy and svm. For each tag, SO provides a textual description, typically consisting of a few sentences. Tags are heavily moderated by humans and automated processes, and are deduplicated through a dictionary of synonyms. We focus on tags that are used at least 1,000 times. We split tags into a two groups: 5,083 general tags and 282 tags that refer to programming languages, defined broadly. For the latter, we merge different versions of the same programming language, converting, for instance python-3.x into the generic tag python. When we define skills, we use all available posts, with answers inheriting the tags of the questions they answer. When we analyze users, we only use their answer posts (ignoring the questions they post) to assess the content of their human capital. To ensure sufficiently accurate descriptions of this human capital, we restrict our analysis to users who post at least 10 answers over the entire period our data cover.
Stack Overflow also provides access to yearly surveys of its users, the Stack Overflow Developer Survey. This provides anonymous, self-reported information on salaries, as well as on tools and languages used that can often be linked to tags in SO. We use the survey data from year 2023, which contains around 12,000 U.S. participants who report salary information.
Finally, we scraped job advertisements from “hnhiring.com”, an index of jobs from the website [^3] https://www.hackernews.com/ Hacker News’ “Who is Hiring?” posts. This yielded a total of 50,998 software related job advertisements. We use approximately 46,000 ads to study skill requirements of jobs and 3,949 ads listing salary information to test the ability of skill requirements to predict wage offers.
2.2 Constructing a taxonomy of software skills
Our goal is to construct a taxonomy of skills related to tasks in software development work. Software development tasks can be viewed as recurring categories of programming problems that developers encounter. The ability to provide advice on how to tackle such problems demonstrates skills and expertise in the area. Therefore, SO, with its vast repository of user-generated questions and answers, offers a unique window onto these real-world challenges and the human capital required to meet them.
In order to organize its community and facilitate information retrieval, the SO platform labels each question with one or more of 66,000 curated tags that indicate the specific tools, frameworks, or concepts involved. Tags are deduplicated and moderated by automated bots and human moderators, ensuring relatively high quality. Before constructing a skill taxonomy, we filter tags in two ways. First, we drop tags appearing fewer than 1000 times. Second, we distinguish tags that refer to programming languages rather than concepts or methods. We drop these tags from our skill taxonomy, because we programming languages are rather tools to implement software skills; the same skill could be applied in different languages. While languages themselves are specialized (Valverde and Solé, 2015; Juhász et al., 2026), they are a different category. In fact, we will later analyze the relationship between skills and languages used.
To identify tags that refer to programming languages, we rely on wiki pages, which are available for most common tags. We first filter the pages to find tags that include “language” or “framework” among their keywords. We then manually review these tags to select those that refer to programming languages, yielding a total of 282 programming languages.
Some programming tags refer to specific versions of a language, such as “python-3.x” or “python-3.7”. In these cases, we merge all tags referring to different versions of the same programming language. Yet, there is some ambiguity as to what counts as a programming language (e.g., web frameworks and developer environments, which are not fully fledged programming languages). We remove such tags when analyzing the relation between skills and languages, yielding 114 unambiguous instances of programming languages.
2.2.1 Identifying software skills
With the processed set of tags in hand, we can now identify canonical software skills in our data set. To do so we express the relation between questions and their tags as a bipartite network with 18,154,593 question nodes that connect to 5,083 tags through 33,378,590 edges. We use these tag–question relationships to group related tags into “canonical” areas of software development expertise (“skills”), applying an SBM (Peixoto, 2017, 2014; Gerlach et al., 2018) to detect communities in the bipartite graph that connects questions to tags (Fig. 1a). SBMs essentially estimate a coarsened version of a network. To do so, they identify communities of nodes and estimate parameters that yield parsimonious descriptions of edge distributions within and between communities. To sharpen the boundaries of inferred communities, we remove $20\%$ of tags that are only weakly connected to the other tags in their community (see Appendix A). This yields 247 different communities. Dropping communities with fewer than 3 tags, we retain 237 communities, connected to 4,054 tags, which we will henceforth will refer to as “skills”.
To generate concise names for each skill community, we label them by an example of a programming skill that illustrates the tags in the community. That is, we name each skill after the task it allows to perform. To produce these labels, we prompt the large language model ChatPGT-4.0 as follows:
This process yields labels that capture the essence of the tags comprising each group (see Appendix B), such as develop NLP models and tools for text analysis, establish secure server access and control, and deploy and secure a scalable web app.
2.3 User skill vectors
Skills are rarely valuable in isolation (Neffke, 2019). Therefore, jobs often require multiple skills. When skills complement each other, they are often required in the same jobs (Anderson, 2017; Alabdulkareem et al., 2018; Stephany and Teutloff, 2024). We investigate this skill bundling by identifying which combinations of skills are frequently mastered by the same SO users. To assess to what extent a user masters a skill, we rely only on the answers they post, not on the questions they raise.
We use the skills (defined from the tasks they allow performing) extracted from the question-tag network to describe the activity of individual SO users as a 237-dimensional vector. Entries of this vector count how many answers the user has provided to questions that are labeled with tags that are associated with each corresponding skill. Specifically, we express the user’s expertise in time period $t$ as:
$\vec{S}_{u}^{t}=\begin{bmatrix}X_{u,1}^{t}\\ ...\\ X_{u,\theta}^{t}\\ ...\\ X_{u,N_{\theta}}^{t}\\ \end{bmatrix}$, (1)
where $N_{\theta}=237$ is the total number of skills in our taxonomy and $X_{u,\theta}^{t}$ the number of answers posted in period $t$ by user $u$ to questions labeled with a tag belonging to the community of tags associated with skill $\theta$. We will interpret this number as the user’s expertise intensity in a skill.
2.4 Skill values
We estimate the value of a skill based on salary information from the Stack Overflow developers survey. Because wages differ widely across countries, we focus on 12,000 US respondents that report a salary in 2023. Respondents list the technologies they use, which often correspond to tags in the SO platform. We use tags that are used in both the survey and the general SO database to find closely matching survey respondents for each SO user. Selecting the 300 most similar survey respondents, we calculate the weighted average salary reported by these respondents. We will use this average as a rough estimate of the wage that the user could command on the US labor market:
where $V_{u}$ is user $u$’s inferred US wage, $V_{\rho}$ the wage of survey respondent $\rho$, and $R_{u}$ the set of survey respondents we matched to user $u$. Weights, $M_{u,\rho}$, are given by the number of tags that user $u$ and respondent $\rho$ share.
To assign a value to a skill, we next calculate the weighted average salary of users possessing the skill, using the skill’s intensity in users’ skill vectors for the years 2018-2023 (the period in which we collect job ads) as weights:
$V_{\theta}$ $=\sum_{u}\frac{X_{u,\theta}^{2018,2023}}{\sum_{u^{\prime}}X_{u^{\prime},\theta}^{2018,2023}}V_{u}.$ (3)
2.5 Constructing a skill space
To determine the relatedness between skills, we analyze which skills are often possessed by the same users. To do so, we collect the skill vectors of all users in an $(N_{\theta}\times N_{u})$ matrix S, with $N_{u}$ the total number of users in the dataset and $N_{\theta}=237$ the total number of skills. Next, we ask how often two skills co-occur within the same users by multiplying $1(\textbf{S}>0)$ with its transpose: $1(\textbf{S}>0)1(\textbf{S}>0)^{\prime}$, where $1(.)$ is an indicator function that evaluates to 1 if its argument is true. Finally, we calculate the Pointwise Mutual Information ($\mathrm{PMI}$) of probability $p_{\theta,\kappa}$ that skill $\theta$ and $\kappa$ co-occur in the same user to capture the amount of (information-theoretic) surprise involved in observing the joint probability $p_{\theta,\kappa}$ against a benchmark where draws of skills $\theta$ and $\kappa$ are independent:
$\mathrm{R}_{\theta,\kappa}=\mathrm{PMI}_{\theta,\kappa}=\log\frac{p_{\theta,\kappa}}{p_{\theta}p_{\kappa}},$ (4)
where $p_{\theta}$ and $p_{\kappa}$ are marginal probabilities, i.e., $p_{\theta}=\sum_{\gamma}p_{\theta,\gamma}$ and $p_{\kappa}=\sum_{\mu}p_{\mu,\kappa}$. We estimate $\mathrm{PMI}$ values and their credible intervals using a Bayesian statistical framework developed by (van Dam et al., 2023). We will refer to these values as the relatedness between skills.
We use these relatedness estimates to construct a skill space, a network where nodes represent skills connected by edges that indicate the relatedness between them. In doing so, we only connect skills with positive relatedness values, ignoring negative relatedness (where skills co-occur less frequently than our random null model predicts) by setting all negative $\mathrm{PMI}$ values to 0. To help interpret this skill space, we embed the relatedness network in a 5-dimensional space using UMAP (McInnes et al., 2018). Next, we cluster sets of contiguous skills using HDBSCAN (McInnes et al., 2017). These clusters are represented in colors of Fig. 1b. The node coordinates in this figure are determined by a UMAP embedding of the relatedness network in a 2-dimensional space. Note, however, that neither the clusters nor the visual representation of the skill space will play a role in our quantitative analyses.
Clusters often group skills with comparable labor market returns. This is illustrated in Fig. 1d, where, we plot the inferred skill values of eq. (3). For example, on the right side of the skill space, we find three clusters revolving around website development that are typically linked to lower wages. By contrast, skills in the top center—focusing on advanced iOS and Android development—offer significantly higher earnings. Additional high-value clusters include those for AI and machine learning (AI/ML), Advanced Programming Concepts, Cloud Computing, and DevOps. The table of Fig. 1e lists the five highest-paying skills, three of which are related to iOS app development, and the five lowest-paying skills, all associated with more basic web development.
Notably, clusters that are conceptually linked—such as AI/ML and Statistics and Data Analysis, and iOS and Android development—tend to lie near each other in the skill space. To further investigate the network structure of the skill space, we can examine individual skills in greater detail. For example, the ten skills most closely related to developing AI models (Fig. 1c) reveal strong connections to other skills in statistics, as well as skills related to image and natural language processing, data visualization, and packaging models into standalone applications. Fig. B.2 shows changes in the relative importance of different skills, highlighting important areas of growth around skills in web frameworks, but also in high-value skills related to AI, cloud computing and Android.
Figure 1: Mapping software skills. a. Stylized depiction of the bipartite question-tag network. SBM groups tags into communities (skills) that connect to similar sets of questions. ChatGPT-4.0 finds a common label that summarizes each community’s tag information. b. Skill space. Pointwise mutual information (relatedness) expresses how surprisingly often two skills are possessed by the same users. UMAP embeds the resulting relatedness network in a 2-dimensional plane (the skill space). c. Close-up on the skill to develop AI models, depicting the original network structure among the 10 most closely related skills. d. Skill values. Nodes are colored according to their skill value, estimated from salary information in the 2023 SO developers survey. Darker shades indicate more valuable skills. e. Table of the five most and least valuable skills.